# references.bib

### v1

@@ IV-B Reinforcement Learning
In order to ensure that the RL policy is able to learn the reward function in a noisy environment a curriculum learning strategy is applied. In order to implement this Equation 20 is adapted to:
${\boldsymbol{x}^{E}_{a}}^{\prime}=\boldsymbol{x}^{E}_{a}+\alpha\mathcal{N}(0,\,\sqrt{diag(\text{CRLB})})$ (28)
where $\alpha$ is a constant that follows a piecewise growth scheduled depicted in Figure 5. By implementing this, the RL policy is able to first converge upon an optimal policy with exact position observations of the cat before having to generalize to incorporate noise.
Figure 5: The piecewise growth schedule of $\alpha$ during training.
During training, the cat is modeled as a chaser spacecraft with no thrust control. The spacecraft then drifts in the vicinity of the mouse, requiring the mouse to adjust its position to maintain a minimum distance. Depending on the initial position and velocity of the chaser spacecraft in the Hill frame, its path can vary drastically. An image depicting several different paths from randomly generated episodes can be seen in Figure 6.
Finally, we use Soft Actor-Critic (SAC) to train our policy which is an off-policy RL algorithm. We chose SAC for its entropy maximization strategy, which is well suited for learning in environments with large amounts of noise.
@@
IV-C MPC control
@@
In equations 29-32 $\mathbb{e}$ is the state error, $Q$ and $R$ are weight matrices, $A$ and $B$ are the constant system and control matrices, and $\mathbb{u}_{lb}$ and $\mathbb{u}_{up}$ are the lower and upper bounds of the thrust control. The system and control matrices are derived using Equations 10-12. The notation $i|k$ implies i steps past k, where k is the current state of the system. During implementation $M$ is set to $8$, thus giving 8 thrust inputs for the mouse. However, only the first computed input correlating to timestep $i=0$ is used. Once this is done, the thrust is applied to the satellite in simulation, and the environment is propagated a single timestep.
V Experiments and Results
In this section we will discuss the implementation of our proposed method, real-world data used for experimentation, and the results of our experiments. During testing, 3 RL models were used – each trained on the same environment with different seeds. A training curve of the RL model can be seen in Figure 7. In both simulation and testing a constellation of 60 satellites was used, and the cat and mouse satellites each had a mass of 2,500kg. Mouse thrust was limited to 1N on each axis. During testing, cat behavior was determined using real-world data of unwarranted fly-bys, and RL performance was compared against a common satellite obstacle avoidance algorithm as well as a greedy optimization algorithm.
V-A Constrained Reinforcement Learning Implementation
At each timestep the EKF, RL policy, and MPC controller are all run in sequential order. In order to ensure the RL policy always performs within acceptable bounds, a constrained RL control scheme is implemented:
$\boldsymbol{a}_{i}=\Biggl{\{}\begin{array}[]{lr}\text{if }\Delta d_{c}<d_{1}&\sim\pi_{\phi}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i})\\ \text{else if }\Delta d_{c}<d_{2}&\boldsymbol{0}\\ \text{else }&-\boldsymbol{x}^{H}_{m,i}\end{array}$ (36)
this restricts the RL policy to computing positional control only when the adversary is predicted to be within a certain range of the mouse spacecraft. If the cat is assumed to be outside of a predetermined distance from the origin, the mouse is provided the necessary control to return to the origin. In addition, if the cat is predicted to be within these two tolerances, no control is provided to the mouse. The constraint policy in Equation 36 is essential for restricting the mouse’s behavior to an acceptable range of possibilities, and provides a high level of assurance during deployment.
V-B Testing Data
For each event, the position of the non-cooperative space craft in relation to the vehicle it approached was computed and used as the cat’s trajectory in simulation. The process of acquiring and processing the data is described in depth in the appendix. An image of each of these relative trajectories can be seen in Figure 8.
V-C Baselines
In order to properly understand the benefits and limitations to deploying a constrained RL policy, we compare the RL policy’s performance against two baselines:
@@ V-C Baselines
  • Delta-V Optimization (DVO): This method, adapted from Slater Et al. [doi:10.2514/1.16812] uses a closed for solution of the Clohesy Whiltshire equations to compute the minimum delta-v required to reach a minimum distance from the origin of the Hill frame. The approach is representative of current industry standards, which generally compute the necessary delta-v to alter a spacecrafts orbit enough that the probability of collision is near zero [BONNAL2020637, Patnala2024], however, is not focused on returning, or minimizing deviation from the initial orbit if possible.
  • Greedy Recursive Search (GRS): In this method we implement a recurseive search algorithm that searches the potential action space of the environment, and computes the position command that will maximize the reward function described in Equation 27 by simulating the performance of the MPC controller. This approach is representative of a greedy optimization algorithm, which in theory should select the optimal solution at each timestep.
Algorithm performance during testing is evaluated using 27, which as discussed prioritizes distance from adversary, deviation from initial orbit, and fuel consumption. An in depth description of how both algorithms are formulated can be found in the appendix. During testing, each method was given full control of the mouse vehicle, while the cat’s trajectory was determined by the obtained TLE data. For each test scenario and method, results were averaged over 3 different seeds tested for 100 runs each.
V-D Results
TABLE I: Test Results
Algorithm Luch/Intelsat 37a Luch2/Eutelsat
DVO $135\pm 30.6$ $143\pm 23.9$
GRS $803\pm 23.2$ $568\pm 34.5$
RL $\boldsymbol{830\pm 15.1}$ $\boldsymbol{607\pm 19.8}$
VI Discussion
The results in the previous section provide considerable insight into the capabilities and advantages to implementing different control schemes for satellite evasion. Possibly the most evident is the limitations of the DVO algorithm, which well optimizes fuel and keeping a minimum distance from the uncooperative spacecraft, however, has no mechanism for maintaining minimum deviation from the initial orbit. This is exacerbated by the orbital dynamics at play: any deviation in the x or y axis of the local frame results in a change in orbit radius which will result in the spacecraft further deviating from its initial orbit with increasing acceleration unless there is some form of intervention. DVO is well representative of many of today’s standards for space-based avoidance which were largely designed for systems that both lack maneuverability and sensing capabilities.
TABLE II: Localization Noise
Constellation Size $\sigma_{X}$ (km) $\sigma_{Y}$ (km) $\sigma_{Z}$ (km)
30 5,542 5,173 2,045
60 4.703 2.446 37.90
100 0.239 0.003 0.147
150 0.193 0.003 0.117
200 0.165 0.004 0.099
In addition to demonstrating limitations of current standards in satellite avoidance, we will highlight some of the key benefits to implementing a Constrained Reinforcement Learning policy for satellite control compared to a more common greedy optimization algorithm. Primarily, how its formulation of the environment as a Markov Decision Process (MDP) allows it to better discover optimal policies in environments with highly complex transition functions. Perhaps this is best seen in Figure 9 when comparing the much smoother trajectories of the RL policy to that of the GRS algorithm. We attribute this to the fact that the GRS algorithm does not take into consideration the potential for maximizing performance given the future state it will be in after taking an action. Consequentially, the algorithm picks actions that do not allow for smooth maneuvering to future positional commands due to the dynamics of the simulation and capabilities of the MPC. This not only results in greater fuel cost, but also slows down the mouse’s ability to travel larger distances in shorter periods of time: providing some insight to its low average deviation from the origin and high number of timesteps spent within $d_{tol}$ of the adversary. It is also worth further discussing that while the RL policy outperformed the GRS algorithm in overall reward, timesteps spent in the vicinity of the cat, and fuel consumed, it did not outperform the GRS algorithm in average deviation from the origin. This may be due to several aspects of the GRS algorithm, including simply choosing optimal control that is not far from the origin, inadequate handling of noisy observation data, and slowed maneuvering due to lack of foresight as previously mentioned.
While we have provided several key benefits and limitations to the different control schemes implemented, we also recognize that the performance of all methods can be improved by reducing the noise of the location estimates provided in the observation. Although the CRLB Bounds provided by Ho et al. give a strict limit on the performance of any TDOA localization algorithm, the accuracy of the algorithm can be improved by increasing the number of RF sensors. For this experiment, we chose a constellation of 60 satellites which results in approximately 10 to 25 LEO satellites being within a $2^{\circ}$ beam transmitted from GEO at any given time. The size of the constellation used was chosen to reflect a feasible number of assets given current constellation sizes as well as a manageable noise level. In table II we have provided the average CRLB bounds of a number of different constellation sizes in order to compare the localization capabilities of the 60-asset constellation to others. In the table, it can be seen that once a constellation approaches a 100 asset size the localization accuracy can be brought down to within 100m, however, this is unrealistic given current satellite constellation magnitudes. A constellation that is too small provides astronomical error which is insufficient for any type of evasion. Finally, a constellation of 60 provides an agreeable middle-ground. Although kilometer accuracy is not ideal we have shown it is sufficient for intelligent evasion. Additionally, it should be noted that the large Error in the Z direction is generally due to outliers which are multiple magnitudes greater than the average noise as shown in Figure 4, and often filtered out. While uncommon these outliers occur due to the configuration of sensors and Dillution of Precission (DOP).
VII Conclusion
Our work proposes an architecture for improved SSA and evasive control by leveraging RF informed localization techniques and a constrained RL control scheme. In this work we have provided the necessary simulation design for training such control schemes, and validated our work by testing on real-world, open data about uncooperative satellite maneuvers. Finally, we compare the RL policy’s performance to an optimization-based algorithm tailored for cat and mouse games, as well as a standard obstacle avoidance algorithm common in satellite avoidance today. We demonstrated that neither alternative outperforms the RL policy on established performance metrics. Our work highlights several key advantages to deploying RL methods in the space environment. First we establish RL’s ability to use noisy and less conventional situational awareness data to establish a robust control policy and second, we establish RL’s ability to learn and navigate the complex dynamics of the space environment. Furthermore, RL provides the additional benefit of being computationally light weight: the model used during experimentation is comprised of 2 layers of 256 nodes, and takes less time to run on a CPU than either optimization algorithm implemented. As the space domain continues to develop and become more congested the current SSA capabilities and planning options will not provide the necessary reactiveness and mobility required of future spacecraft. Space-based RF sensing unlocks a new frontier in satellite autonomy capabilities; when integrated with advanced control algorithms—such as reinforcement learning-based policies—it has the potential to expand the boundaries of autonomous satellite maneuvering.
Acknowledgments
Appendix A Extended Kalman Filter
As discussed in the Methods section, we employ an Extended Kalman Filter (EKF) to filter out some of the extreme noise found in the TDOA location estimates. The main update function has been described in Equation 21, however, there are several other steps during implementation not yet discussed. At each timestep, the EKF must start with a reference state vector from the previous timestep $\boldsymbol{x}(t_{i-t})$, and covariance matrix $\boldsymbol{P}_{i-1}$. For the following equations governing the EKF, the state vector $\boldsymbol{x}$ is defined as:
$\boldsymbol{x}_{i}^{-}=\boldsymbol{x}(\boldsymbol{x}(t_{i-t}))$ (38)
Where the function $\boldsymbol{x()}$ represents Cowell’s Formulation, which propagates the state forward a single timestep using Newton’s Equations of motion as discussed in the Methods section. And, $\boldsymbol{\Phi}$ is the state transition matrix computed from a second order Taylor expansion series:
where $\boldsymbol{F}$ is the state matrix of the 2-body problem $\frac{\partial\dot{\boldsymbol{X}}}{\partial\boldsymbol{X}}$. The Kalman gain $\boldsymbol{K}$ is then updated:
This equation is nearly identical to Equation 21 in the Methods section, however, we show its full form here where $\boldsymbol{H}_{i}$ is not taken out. Finally, the covariance matrix is then updated in preparation for the next update:
The updated state vector ${\boldsymbol{x}^{E}_{a,i}}^{+}$ and covariance matrix $\boldsymbol{P}_{i}^{+}$ are then used as the state vector and covariance matrix from the previous timestep in the next update.
During initialization, $\boldsymbol{P}$ is set to the diagnol matrix of the square of the average error standard deviations of the constellation’s position estimates found in Table II, and $R$ was tuned through trial and error to provide the most accurate results.
Appendix B Greedy Recursive Search Algorithm
Our GRS algorithm solves for the optimal position command to provide to the satellite by iterating through an increasingly more granular state space of potential positions within a given tolerance of the averaged cat position estimate. The full algorithm can be seen in Algorithm 1. The algorithm takes the average location of the cat’s most recent position estimates, and compares the potential reward for a range of points that are distance $d_{m}$ from the mean position. $d_{m}$ was chosen by analytically experimenting with values greater than $d_{c}$ in order to allow preemptive evasion. In order to solve for the optimal position command, the function is run recursively to analyze a smaller subset of potential points outside of the optimal point found from the previous iteration. Additionally, if the position of the cat is outside of $d_{2}$, the mouse will return to the origin.
@@ Appendix B Greedy Recursive Search Algorithm
10:              $\boldsymbol{g}^{H}\leftarrow\frac{\sum_{j=i-N}^{i}{\boldsymbol{x}^{H}_{a,j}}^{+}}{N}+\begin{bmatrix}x&y&z\end{bmatrix}$
@@
Appendix C Delta-V Optimization Algorithm
We have adapted a classic delta-v optimization algorithm from the works of Slater et al. for the 3D Hill frame. The algorithm aims to minimize the cost function:
$P=I-\boldsymbol{e}\boldsymbol{e}^{T}$ (45)
where $\boldsymbol{e}$ is a unit vector in the direction of the velocity in the Hill frame after correction. $\phi_{12}$ is the closed form solution for $\boldsymbol{x}^{H}$ using the Clohessy-Whiltshire equations for some fixed time $t$ assuming a starting position at the origin:
$\mid\Delta\boldsymbol{V}\mid=\sqrt{\frac{D^{2}}{\lambda_{max}}}$ (47)
Appendix D Real-World Data Acquisition

### v2

@@ V-A Reinforcement Learning Training Scheme
As discussed in Section III, the noisy position estimates of the cat prevents the policy from knowing the true state of the environment at any given time. And, while this can be used to improve generalization of the policy during testing, it can also hinder the chance of learning an optimal policy. In order to mitigate this a curriculum learning strategy is applied to the localization noise during training, similar to many curriculum-based domain randomization techniques [DBLP:journals/corr/abs-1910-07113, 10606428]. In order to implement this Equation 25 is adapted to:
${\boldsymbol{z}^{E}_{a}}=\boldsymbol{x}^{E}_{a}+\alpha\mathcal{N}(0,\,\sqrt{diag(\text{CRLB})})$ (30)
where $\alpha$ is a constant that follows a piecewise growth scheduled. By implementing this, the RL policy is able to first converge upon an optimal policy with exact position observations of the cat before having to generalize to incorporate noise.
Additionally, it is important during training that the adversary’s actions represent a sufficiently diverse set of potential behaviors so that the trained policy is able to demonstrate robust and generalizable capabilities during testing. In order to ensure this the cat is modeled as a chaser spacecraft with no thrust control. The spacecraft then drifts in the vicinity of the mouse, requiring the mouse to adjust its position to maintain a minimum distance. Depending on the initial position and velocity of the chaser spacecraft in the Hill frame, its path can vary drastically. An image depicting several different paths from randomly generated episodes can be seen in Figure 5.
@@
V-B Constrained Reinforcement Learning
It is essential that during testing the actions provided by the RL policy always performs within acceptable bounds. Thus, we constrain the potential action set during testing by weighting the probability of taking specific actions. In order to do this, we have identified three potential action scenarios for the cat and mouse simulation: first, the cat is within a predetermined distance tolerance of the mouse and evasive maneuvers should be taken ($d_{c}<c_{1}$), second, the cat is outside of the first distance tolerance but not sufficiently far and thus best to wait to understand its behavior ($c_{1}<d_{c}<c_{2}$), and third, the cat is sufficiently far from the mouse and the spacecraft should return to its initial orbit ($d_{c}>c_{2}$). This allows us to define the potential action space during testing:
$\boldsymbol{A}_{i}=\Biggl{\{}\begin{array}[]{lr}\sim\pi_{\phi}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i})\\ \boldsymbol{0}\\ -\boldsymbol{x}^{H}_{m,i}\end{array}$ (34)
where $\sim\pi_{\phi}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i})$ is the action sampled from the RL policy, $\boldsymbol{0}$ would entail no change in the mouse’s position, and $-\boldsymbol{x}^{H}_{m,i}$ is the necessary change in position to return to the origin. Each action represents the scenarios mentioned in their respective order. At each timestep during testing actions are sampled following discrete probability distribution which represents the probability of each of the three previously defined inequalities being true ($\pi_{s1},\pi_{s2},\pi_{s3}$).
Each probability is computed by taking the weighted sum of the probability of each inequality for the past $N$ estimates of the cat’s position:
${\pi_{s}}_{k}=\sum_{j=1}^{N}w_{j}P(b\mid{{\boldsymbol{z}_{a,i-N}^{H}}^{+}},\boldsymbol{x}_{m,i}^{H})$ (35)
where $b$ represents one of the three potential inequalities for each scenario, and $w$ comes from a set of normalized weights. The function $P(b\mid{{\boldsymbol{z}_{a,i-N}^{H}}^{+}},\boldsymbol{x}_{m,i}^{H})$ is the cumulative density function (CDF) of the euclidean distance between the cat and mouse given an observed location. This can be computed by treating it as a noncentral chi-squared distribution with three dimensions:
$P(b\mid\boldsymbol{x})=1$ (36)
$-\left[Q(\sqrt{\hat{c}}-\sqrt{\lambda})+Q(\sqrt{\hat{c}}+\sqrt{\lambda})+\sqrt{\frac{2}{\pi}}\frac{sinh(\sqrt{\lambda\hat{c}})}{\sqrt{\lambda}}e^{-\frac{\hat{c}+\lambda}{2}}\right]$
where $\boldsymbol{x}$ is the mean of the point of interest equal to ${{\boldsymbol{z}_{a,i-N}^{H}}^{+}}-\boldsymbol{x}_{m,i}^{H}$, $Q()$ is a gaussian Q function:
[…]
$\hat{c}$ is the square of the distance tolerance normalized by the standard deviation $\sigma$ of the position estimate:
$\hat{c}=\left(\frac{c}{\sigma}\right)^{2}$ (38)
and lambda is the sum of the squares of the position estimate again normalized by the standard deviation:
$\lambda(\boldsymbol{x})=\sum_{j=1}^{3}\left(\frac{x_{j}}{\sigma}\right)^{2}$ (39)
Figure 6: The CDF function for each of the three defined inequalities.
A plot of the CDF function for each of the three inequalities represented by $b$ can be seen in Figure 6. The noncentral chi-squared distribution models the distribution of the sum of square of k independent, standard normally distributed variables. This is equal to the square of the euclidean distance of a noisy position estimate assuming that there is uniform standard deviation ($\sigma$) of variables across all axis. During testing, $\sigma$ can be computed by taking the standard deviation of the distance between the cat’s position estimate and mouse’s location over the last $N$ estimates.
V-C MPC control
@@
In equations 40-43 $\mathbb{e}$ is the state error, $Q$ and $R$ are weight matrices, $A$ and $B$ are the constant system and control matrices, and $\mathbb{u}_{lb}$ and $\mathbb{u}_{up}$ are the lower and upper bounds of the thrust control. The system and control matrices are derived using Equations 15-17. The notation $i|k$ implies i steps past k, where k is the current state of the system. During implementation $M$ is set to $8$, thus giving 8 thrust inputs for the mouse. However, only the first computed input correlating to timestep $i=0$ is used. Once this is done, the thrust is applied to the satellite in simulation, and the environment is propagated a single timestep.
VI Baselines
We test our method against two deferent baselines chosen to represent the different methodologies currently seen in satellite collision avoidance or evasion. The first is a single burn delta-v maneuver designed to ensure the spacecraft misses a potential obstacle by a minimum distance. The second is an optimization based algorithm that attempts to solve for the optimal position at each timestep with respect to the same reward function used to train the RL policy. In addition, because these methods are not designed for extremely noisy data we find it beneficial to pass all observations through an Extended Kalman Filter (EKF) before being fed to either algorithm. Observation variables that have been filtered by the EKF are notated by the superscript $+$. A more detailed explanation of the EKF process can be found in Appendix A.
VI-A Delta-V Optimization
@@
The first baseline implemented is a Delta-V Optimization method which will be refereed to as DVO in further sections. This method, adapted from Slater Et al. [doi:10.2514/1.16812] uses a closed for solution of the Clohesy Whiltshire equations to compute the minimum delta-v required to reach a minimum distance from the origin of the Hill frame. This approach is representative of current industry standards, which generally compute the necessary delta-v to alter a spacecrafts orbit enough that the probability of collision is near zero [BONNAL2020637, Patnala2024]. The algorithm aims to minimize the cost function:
$P=I-\boldsymbol{e}\boldsymbol{e}^{T}$ (45)
where $\boldsymbol{e}$ is a unit vector in the direction of the velocity in the Hill frame after correction. $\phi_{12}$ is the closed form solution of the Clohessy-Whiltshire equations for some fixed time $t$ assuming a starting position at the origin:
$\mid\Delta\boldsymbol{V}\mid=\sqrt{\frac{D^{2}}{\lambda_{max}}}$ (47)
VI-B Greedy Recursive Search
The DVO algorithm well characterizes the majority of satellite avoidance methods currently utilized today. However, it was not necessarily designed to optimize the cat and mouse problem we are exploring, and thus we provide a second optimization based method more tailored to the problem at hand known as Greedy Recursive Search (GRS). Our GRS algorithm solves for the optimal position command to provide to the satellite by iterating through an increasingly more granular state space of potential positions within a given tolerance of the averaged cat position estimate. The algorithm attempts to optimize the same reward function used to train the RL policy described in Equation 29, and the full algorithm can be seen in Algorithm 1.
@@ VI-B Greedy Recursive Search
10:              $\boldsymbol{g}^{H}\leftarrow\frac{\sum_{j=i-N}^{i}{\boldsymbol{z}^{H}_{a,j}}^{+}}{N}+\begin{bmatrix}x&y&z\end{bmatrix}$
@@
In Algorithm 1 $\phi_{min},\phi_{max},\theta_{min},$ and $\theta_{max}$ represent angle ranges in polar coordinates, they are generally set to 00 and $2\pi$ when the algorithm is called. The algorithm takes the average location of the cat’s most recent position estimates, and compares the potential reward for a range of points that are distance $d_{m}$ from the mean position. $d_{m}$ was chosen by analytically experimenting with values greater than $c_{1}$ in order to allow preemptive evasion. In order to solve for the optimal position command, the function is run recursively to analyze a smaller subset of potential points outside of the optimal point found from the previous iteration. Additionally, if the position of the cat is outside of $c_{2}$, the mouse will return to the origin like the action space defined for our constrained RL policy.
VII Experiments and Results
In this section we will discuss the implementation of our proposed method, real-world data used for experimentation, and the results of our experiments. We trained our policy using Soft Actor-Critic (SAC) which is an off-policy RL algorithm. We chose SAC for its entropy maximization strategy, which is well suited for learning in environments with large amounts of noise. During testing, 3 RL models were used – each trained on the same environment with different seeds. A training curve of the RL model can be seen in Figure 7. In both simulation and testing a constellation of 60 satellites was used, and the cat and mouse satellites each had a mass of 2,500kg. Mouse thrust was limited to 1N on each axis. During testing, cat behavior was determined using real-world data of unwarranted fly-bys, and RL performance was compared against a common satellite obstacle avoidance algorithm as well as a greedy optimization algorithm.
VII-A Testing Data
For each event, the position of the non-cooperative space craft in relation to the vehicle it approached was computed and used as the cat’s trajectory in simulation. The process of acquiring and processing the data is described in depth in Appendix 2. An image of each of these relative trajectories can be seen in Figure 8.
Both our method’s and the implemented baselines’ performance during testing is evaluated using 29, which as discussed prioritizes distance from adversary, deviation from initial orbit, and fuel consumption. During testing, each method was given full control of the mouse vehicle, while the cat’s trajectory was determined by the obtained TLE data. For each test scenario and method, results were averaged over 3 different seeds tested for 100 runs each.
VII-B Results
TABLE I: Test Results
Algorithm Luch/Intelsat 37a Luch2/Eutelsat
DVO $135\pm 30.6$ $143\pm 23.9$
GRS $807\pm 23.3$ $568\pm 34.5$
RL $\boldsymbol{821\pm 33.4}$ $\boldsymbol{610\pm 38.2}$
VIII Discussion
The results in the previous section provide considerable insight into the capabilities and advantages of implementing different control schemes for satellite evasion. Most evident are the limitations of the DVO algorithm, which optimizes fuel and maintains a minimum distance from the uncooperative spacecraft, however, has no mechanism for maintaining minimum deviation from the initial orbit. This is exacerbated by the orbital dynamics: any deviation in the x or y axis of the local frame results in a change in orbit radius which will result in the spacecraft further deviating from its initial orbit with increasing acceleration unless there is some form of intervention. DVO is representative of many of today’s standards for space-based avoidance, which were largely designed for systems that both lack maneuverability and sensing capabilities.
TABLE II: Localization Noise
Constellation Size $\sigma_{X}$ (km) $\sigma_{Y}$ (km) $\sigma_{Z}$ (km)
30 5,542 5,173 2,045
60 4.703 2.446 37.90
100 0.239 0.003 0.147
150 0.193 0.003 0.117
200 0.165 0.004 0.099
In addition to demonstrating limitations of current standards in satellite avoidance, we will highlight some of the key benefits to implementing a constrained Reinforcement Learning policy for satellite control compared to a more common greedy optimization algorithm. Primarily, how its formulation of the environment as an MDP allows it to better optimize control over a longer horizon. Perhaps this is best seen in Figure 9 when comparing the much smoother trajectories of the RL policy to that of the GRS algorithm. We attribute this to the fact that the GRS algorithm does not take into consideration the potential for maximizing performance given its future state after taking an action. Consequentially, the algorithm picks actions that do not allow for smooth maneuvering to future positional commands due to the dynamics of the simulation and capabilities of the MPC. This not only results in greater fuel cost, but also slows down the mouse’s ability to travel larger distances in shorter periods of time: providing some insight to its low average deviation from the origin and high number of timesteps spent within $d_{tol}$ of the adversary. Additionally notable, while the RL policy outperformed the GRS algorithm in overall reward, timesteps spent in the vicinity of the cat, and fuel consumed, it did not outperform the GRS algorithm in average deviation from the origin. This may be due to several aspects of the GRS algorithm, including simply choosing optimal control that is not far from the origin, inadequate handling of noisy observation data, and slowed maneuvering due to shorter horizon planning as previously mentioned.
While we have provided several key benefits and limitations to the different control schemes implemented, we also recognize that the performance of all methods can be improved by reducing the noise of the location estimates provided in the observation. Although the CRLB Bounds provided by Ho et al. give a strict limit on the performance of any TDOA localization algorithm, the accuracy of the algorithm can be improved by increasing the number of RF sensors. For this experiment, we chose a constellation of 60 satellites which results in approximately 10 to 25 LEO satellites within a $2^{\circ}$ beam transmitted from GEO at any given time. The size of the constellation was selected to reflect a feasible number of assets given current constellation sizes as well as a manageable noise level. In table II we have provided the average CRLB bounds of a number of different constellation sizes in order to compare the localization capabilities of the 60-asset constellation to others. Once a constellation approaches 100 assets the localization accuracy can be brought down to within 100m, however, this is unrealistic given current satellite constellation magnitudes. A constellation that is too small provides error, which is insufficient for any type of evasion. Finally, a constellation of 60 provides a sufficient tradeoff. While more finite accuracies are more desirable, we have shown it is sufficient for intelligent evasion. Additionally, it should be noted that the large Error in the Z direction is generally due to outliers which are multiple magnitudes greater than the average noise as shown in Figure 4. While uncommon these outliers occur due to the configuration of sensors and Dilution of Precision (DOP).
IX Conclusion
Our work proposes an architecture for improved SDA and evasive control by leveraging RF informed localization techniques and a constrained RL control scheme. In this work we have provided the necessary simulation design for training such control schemes, and validated our work by testing on real-world, open data about uncooperative satellite maneuvers. Finally, we compare the constrained RL policy’s performance to an optimization-based algorithm tailored for cat and mouse games, as well as a standard obstacle avoidance algorithm common in satellite avoidance today. We demonstrated that neither alternative outperforms the RL policy on established performance metrics. Our work highlights several key advantages to deploying RL methods in the space environment. First we establish RL’s ability to use noisy, but ubiquitous RF data to establish a robust control policy and second, we establish RL’s ability to learn and navigate the complex dynamics of the space environment. Furthermore, RL provides the additional benefit of being computationally light weight: the model used during experimentation is comprised of 2 layers of 256 nodes, and takes less time to run on a CPU than either optimization algorithm implemented. As the space domain continues to develop and become more congested the current SDA capabilities and planning options will not provide the necessary reactivity and mobility required of future spacecraft. Space-based RF sensing unlocks a new frontier in satellite autonomy capabilities; when integrated with advanced control algorithms — such as RL-based policies — it has the potential to expand the boundaries and improve the accessibility of autonomous satellite maneuvering.
Acknowledgments
Appendix A Extended Kalman Filter
As mentioned in Section VI, the position estimates provided by the TDOA localization algorithm are extremely noisy. In order to mitigate errors in our baselines caused by this an EKF is implemented in order to smooth the position data before it is fed as intput to either of the baseline control algorithms. The EKF implemented follows an architecture similar to that often seen in space systems [springerSatelliteOrbits]. At each timestep, the EKF must start with a reference state vector from the previous timestep $\boldsymbol{x}(t_{i-t})$, and covariance matrix $\boldsymbol{P}_{i-1}$. For the following equations governing the EKF, the state vector $\boldsymbol{x}$ is defined as:
$\boldsymbol{x}_{i}^{-}=\boldsymbol{x}(\boldsymbol{x}(t_{i-t}))$ (49)
Where the function $\boldsymbol{x()}$ represents Cowell’s Formulation, which propagates the state forward a single timestep using Newton’s Equations of motion:
$\boldsymbol{F}=\frac{Gm_{1}m_{2}}{r^{3}}\hat{\boldsymbol{r}}$ (51)
And, $\boldsymbol{\Phi}$ is the state transition matrix computed from a second order Taylor expansion series:
where $\boldsymbol{F}$ is the state matrix of the 2-body problem $\frac{\partial\dot{\boldsymbol{X}}}{\partial\boldsymbol{X}}$ derived from equation 51. The Kalman gain $\boldsymbol{K}$ is then updated:
This equation is nearly identical to Equation LABEL:eq:ekf1 in the Methods section, however, we show its full form here where $\boldsymbol{H}_{i}$ is not taken out. Finally, the covariance matrix is then updated in preparation for the next update:
The updated state vector ${\boldsymbol{x}^{E}_{a,i}}^{+}$ and covariance matrix $\boldsymbol{P}_{i}^{+}$ are then used as the state vector and covariance matrix from the previous timestep in the next update. The position of the updated state vector ${\boldsymbol{x}^{E}_{a,i}}^{+}$ is then translated into the Hill frame following Equation 1 in order to provide ${\boldsymbol{x}^{H}_{a,i}}^{+}$ to the control algorithm.
Additionally, during initialization $\boldsymbol{P}$ is a diagonal matrix of the square of the average error standard deviations of the constellation’s position estimates found in Table II, and $R$ was tuned through trial and error to provide the most accurate results.
Appendix B Real-World Data Acquisition